MAT 123, INTRODUCTION TO CALCULUS, SUNY AT STONY BROOK, Spring 2003

Text: Selected Chapters from PreCaluclus and Single Variable Calculus, James Steward.

Syllabus: We will cover approximately Chapters 2,3,4,5 and 7 in the precalculus section and Chapter 2 in the calculus part, with omissions of certain sections. A more detailed syllabus will be provided.

Classdays. Classes begin on Wednesday, January 22 and end on Friday, May 9. Holidays: There are no classes during spring break, March 17-23. There are also no

classes during April 16-18, for the observance of Passover.

Lecture times and locations:

day  time  room  instructor
LEC 1 MF 12:40pm- 2:00pm HARRIMAN HLL 137 Bishop,Christopher
R01  M 10:30am-11:25am SOCBEHAV SCI N436 Baum, Kevin
R02 Th 9:50am-10:45am EARTH&SPACE 79 Sugrim, James
R03  Tu 12:50pm- 1:45pm MELVILLE LBR N3063 Lamb, Corrinne
R04  Tu 9:50am-10:45am HARRIMAN HLL 115 Baum, Kevin
R05 W 9:25am-10:20am EARTH&SPACE 69 Shah, Samir
R06  Th 12:50pm- 1:45pm MELVILLE LBR N3063 Lamb, Corrinne
LEC 2 TuTh 3:50pm- 5:10pm EARTH&SPACE 1  Bernhard,William S.
R07  W 10:30am-11:25am MELVILLE LBR N4000 Shah, Samir
R08  M  4:25pm- 5:20pm OLD CHEM 138  Chen, Xiaojun
R09  W  4:25pm- 5:20pm OLD CHEM 138  Chen, Xiaojun
R10  F 10:30am-11:25am MELVILLE LBR N4000 Sharma, Gaurav
R11  M  9:25am-10:20am OLD CHEM 138  Wong,Wai-Ho (Patrick)
R12  Th 11:20am-12:15pm PHYSICS P117  Sugrim, James
R13  F  9:25am-10:20am OLD CHEM 138  Sharma, Gaurav
ELC 3 TuTh 7:00pm- 8:20pm EARTH&SPACE 131  Ambrose, Donald
ELC 4 MW 7:00pm- 8:20pm HARRIMAN HLL 112 Kumpel, Paul

There is no recitation for Lectures 3 and 4. The grader for both of these lectures is Praveen Ramachandran.

Instructor  Office Office hours  email
Bishop  4-112 MW 10:30-11:30am  bishop@math.sunysb.edu
Bernhard  4-110 M 4:30-5:30 in MLC  bill@math.sunysb.edu

Ambrose  4-110 TuTh 6-7pm  dambrose@ic.sunysb.edu
Kumpel  3-120  pkumpel@math.sunysb.edu
Baum  Th 10-12  lax11570@hotmail.com
Sugrim  Tu 5:00-7:00pm in MLC ssugrim@mathlab.sunysb.edu

Lamb  TT 1:45-2:15 N3063 Lib, clamb@ic.sunysb.edu

W 1-2 MLC

Shah  bluepapaya@hotmail.com
Chen  W 7:00-9:00pm  chen@math.sunysb.edu
Sharma  2-105  gsharma@ic.sunysb.edu
Wong  M 10:30-12:30  PWong312@aol.com
Ramachandran  praveen_ramachandran81@yahoo.com

Homework. A list of problems from the text will be handed out. The recitation instructor will assign, collect and return homework. In general, problems are due the week after the section has been covered in lecture.

Examinations. There will be two midterm tests: Monday, March 3 and Wednesday, April 9. These will be given at night (8:30pm-10:00pm) in a location which will be annouced in class. The final exam is on Thursday, May 15 at 11:00-1:30. The location will be announced in class during the last week of the semester. No calculators, notes or books will be allowed during exams. Bring your student ID to all exams.

Grades. Grades are computed using: Recitation 10%, Midterms 25% each, Final 40%. Disabilities. If you have any condition, such as physical or mental disability, that will

make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: This is in a new location: S-240A in the basement of the math building. MLC provides help with pre-calculus and calculus material. It is staffed most days and some evenings; an exact schedule will be posted outside the room.

Drop dates: Last day for Add/Drop is January 28. Last day to switch to lower numbered math course is Friday, March 7. The last day to withdraw from a course is March 28.